Determine the monthly payment if you take out a loan for $ 230000 with an annual interest rate of 7% and the term of the loan is 20 years.
Monthly Payment =

Answer by Theo(13342)

loan is for 230,000
annual interest rate is 7%.
term of the loan is 20 years.
20 years * 12 = 240 months
7% per year = (7/12)% per month.

using the texas instruments BA-II Plua financial calculator, my inputs are:
pv = 230000
fv = 0 (future value)
n = 240 months (number of time periods)
i/y = 7/12 percent per month (interest rate percent per month)
payments are made at the end of each month.
i click on pmt and it tells me that the payment at the end of each month needs to be 1783.19 rounded to the nearest penny.
it shows the payment as negative because it is money going out, while it shows the present value as positive because it is money coming in.
that convention is required by the calculator.

you can do the same thing using the online financial calculator at https://arachnoid.com/finance/

here are the results from using that calculator.



note that (7/12)% is equal to .58333333.....%
with the ti calculator, you can enter 7/12.
with the online calculator, you have to do the arithmetic and enter the result of 7/12.

both calculators will give you the same answer if you do the inputs correctly.